Find the radius of circle passing through (0,0), (1,0), (0,1)
 

Grade:11

1 Answers

Arun
4 years ago
Dear student
 
As this form a right angled triangle and if we want these points as points on circumference hence (1,0) and k0,1) acts as a diameter end points.
 
Hence diameter = sqrt (2)
 
Hence radius = sqrt(2) / 2 = 1/sqrt(2)